OpenStudy (anonymous):

Find the indicated nth partial sum of the arithmetic sequence. 1.4, 4.6, 7.8, 11, ..., n = 100 A. 7250 B. 16300 C.15980 D.15981.8 E. 15978.2

OpenStudy (anonymous):

Hi, The common difference is (4.6-1.4)=3.2. The 100th term is 1.4+(100-1)*3.2=318.2. The sum of the first 100 terms is then 100/2*(1.4+318.2)=15980.

OpenStudy (anonymous):

Oh ok thank you very much!

OpenStudy (anonymous):

You're welcome!